OpenStudy (anonymous):

What is the simplified mixed number form of forty-five over eighteen? two and one over three two and one over two two and nine over eighteen two and two over three

OpenStudy (anonymous):

how many times does 18 go into 45?

OpenStudy (anonymous):

This is what you're trying to solve \[\frac{ 45 }{ }\]

OpenStudy (anonymous):

\[\frac{ 45 }{ 8 }\]

OpenStudy (anonymous):

45/18=5/2 by dividing each by 9